Seven students went on a diet in an attempt to lose weight, with two of them losing weight while all the others added weight. Is the diet is an effective way to losing weight α=1% (5mks)

1
Expert's answer
2022-03-13T18:51:54-0400

Let p denote the proportion of people who lose weight with the diet. The null hypothesis is that the diet is not effective, i.e. the proportion is 0.5, the alternative is that the proportion is greater than 0.5.

 

We have the

sample proportion

"\hat{p}=\frac{2}{7}"

The P- value

"P\left( \hat{p}\geqslant \frac{2}{7} \right) =P\left( \sqrt{n}\frac{\hat{p}-p}{\sqrt{p\left( 1-p \right)}}\geqslant \sqrt{n}\frac{\frac{2}{7}-p}{\sqrt{p\left( 1-p \right)}} \right) =\\=P\left( Z\geqslant \sqrt{7}\frac{\frac{2}{7}-0.5}{\sqrt{0.5\left( 1-0.5 \right)}} \right) =\\=P\left( Z\geqslant -1.13389 \right) =1-\Phi \left( -1.13389 \right) =0.87158"

Since the P-value is greater than "\alpha =0.01" the null hypothesis is not rejected. The diet is not effective.